hellop plese help me with this. Thankyou.

Two numbers differ by 3. The sum of the larger and onefourth the smaller is 13. What are the numbers??

Answer by stanbon(48558)

Two numbers differ by 3. The sum of the larger and one fourth the smaller is 13. What are the numbers??
-------------
Let the larger be "x+3"
The smaller is "x"
EQUATION:
x+3 + (1/4)x = 13
Add (4/4)x and (1/4)x to get:
(5/4)x = 10
x = 4*10/5
x = 8 (smaller number)
x+3 = 11 (larger number)
============
Cheers,
Stan H.